OpenStudy (mtalhahassan2):

the orbital velocity of a planet must increase if the orbital radius decreases

OpenStudy (mtalhahassan2):

true/ false

OpenStudy (mtalhahassan2):

@Miracrown

OpenStudy (mtalhahassan2):

@alexis42

OpenStudy (korosh23):



OpenStudy (korosh23):

\[V= \frac{ 2\pi r }{ T }\]

OpenStudy (korosh23):

You can understand the relationship of V and radius if radius increase, velocity increase. If velocity increase, radius increase VT/2pi= r

OpenStudy (korosh23):

Hope this helped :)

OpenStudy (mtalhahassan2):

@korosh23 so its true

OpenStudy (korosh23):

In my opinion yes, but do more research

Parth (parthkohli):

\[\frac{mv^2}r = \frac{GMm}{r^2}\]\[\Rightarrow v^2 = \frac{GM}{r}\]This is one way to look at why the statement is true.

OpenStudy (mtalhahassan2):

so its look true to me

OpenStudy (korosh23):

but in that situation, the new formula, by increasing the radius, you decrease the velocity

OpenStudy (korosh23):

so yes, if radius decreases, that means velocity increases

OpenStudy (korosh23):

Use my formula if you are dealing with Uniform circular motion Use his formula if you are dealing with planets and orbits

OpenStudy (mtalhahassan2):

actually i am dealing with orbits so

OpenStudy (korosh23):

Then use his formula, but it is good to know both situations :)

OpenStudy (korosh23):

Sorry if I confused you at first

OpenStudy (mtalhahassan2):

ok thnx so it be true

OpenStudy (anonymous):

k